OpenStudy (anonymous):

Write an equation in point-slope form of the line that has a slope of 1/2 and passes through the point (5, -2).

OpenStudy (mathmale):

First, please type in the "point-slope form" of the equation of a straight line. Once you have that form, the rest of the work here is quite simple.

OpenStudy (texaschic101):

hello peter...lol. point-slope form is : y - y1 = m(x - x1) slope(m) = 1/2 (5,-2) x1 = 5 and y1 = -2 now just sub, but do not solve, leave the parenthesis in.

OpenStudy (anonymous):

Hi Texas Chic (#101)

OpenStudy (texaschic101):

y - y1 = m(x - x1) y - (-2) = 1/2(x - 5) y + 2 = 1/2(x - 5) ===> here you go :)

OpenStudy (anonymous):

Ohh duh I thought you meant subtract this line (5,-2) x1 = 5 and y1 = -2 I was like what? Thanx

OpenStudy (texaschic101):

lol....no problem :) glad to help